OpenStudy (anonymous):

The number of people in a town of 10,000 who have heard a rumor started by a small group of people is given by the following function: N(t) = 10,000/(5 + 1245e-^.97t) How long before 1000 of the people in the town have heard the rumor?

OpenStudy (anonymous):

man they really reach for these world problems, don't they?

OpenStudy (anonymous):

they really do, it confuses me every time

OpenStudy (anonymous):

@satellite73 do you know how i would begin to solve this?

OpenStudy (anonymous):

yes, set the expression equal to 1000 and solve for \(x\)

OpenStudy (anonymous):

set this exual to 1000? N(t) = 10,000/(5 + 1245e-^.97t)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{10,000}{5+1245e^{-.97t}}=1000\] is the first step

OpenStudy (anonymous):

and solve for t right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

maybe next step would be \[10=5+1245e^{-.97t}\]

OpenStudy (anonymous):

okay thanks, so after that step i would solve for t? or before?

OpenStudy (anonymous):

i got t=0.0106

OpenStudy (anonymous):

that seems unlikely

OpenStudy (anonymous):

im not sure i used a calculator

OpenStudy (anonymous):

starting with \[10=5+1245e^{-.97t}\] what did you do next?

OpenStudy (anonymous):

i put it into a calculator and solved for t and it gave me 0.0106

OpenStudy (anonymous):

yes, i think there must be some mistake here

OpenStudy (anonymous):

yes, thank you so much for your help anyways. i can wait and ask the teacher because they left me without any formulas or anything

OpenStudy (anonymous):

lets make sure it is right, it is not that hard

OpenStudy (anonymous):

\[10=5+1245e^{-.97t}\] subtract 5\[5=1245e^{-.97t}\] divide \[\frac{5}{1245}=\frac{1}{249}=e^{-97t}\] \[-.97t=\ln(\frac{1}{249})=-\ln(249)\] \[t=\frac{\ln(249)}{.97}\]

OpenStudy (anonymous):

thank you again, i understand it much better now